A rectangular painting has length 2 inches less then three times the width. The area is 40 inches squared?
Answer by Boreal(15235)

width=x
length=3x-2
area is product and is 40 in^2
x*(3x-2)=40
3x^2-2x=40
3x^2-2x-40=0
(3x+10)(x-4)=0
only the root x=4 makes sense.
width is 4 in
length is 10 in
area is 40 in^2.